The piston diameter of a certain hand pump is

0.6

inch. The manager determines that the diameters are normally​ distributed, with a standard deviation of

0.009

inch. After recalibrating the production​ machine, the manager randomly selects

20

pistons and determines that the standard deviation is

0.0075

inch. Is there sufficient evidence for the manager to conclude that the standard deviation has decreased at the

α=0.05

level of​ significance?
Answer #1

Claim: The manager to conclude that the standard deviation has decreased.

Sample size = n = 20

Sample standard deviation = s = 0.0075

The null and alternative hypothesis is

Level of significance = α = 0.05

Test statistic is

Degrees of freedom = n - 1 = 20 - 1 = 19

P-value = P( ) = 0.8285

P-value > 0.05 we fail to reject null hypothesis.

Conclusion: There is not sufficient evidence for the manager to conclude that the standard deviation has decreased.